# Post-reduction inference for confidence sets of models

### v1

@@ 5.2. Coverage and expected size of the confidence sets
In the Gaussian context, because the $F$ test is exact, the coverage probability is only affected by the reduced survival probability in the first stage and not by the small sample size at the second stage, and the efficacy of sample splitting is relatively high, the directionality of the $F$ test compensating to a large extent for the smaller sample size, provided that all signal variables survive reduction. In broad terms, therefore, the main aspect influencing the relative advantage of the co-sufficient and ancillary tests over the split-sample tests is the difficulty of the reduction phase. In non-linear regression models, where a split likelihood ratio test has to be used, the coverage probability of $\mathcal{M}$ is also affected by the smaller sample size in the second stage, due to the $\chi^{2}$ approximation to the likelihood-ratio statistic, although this too could be improved by higher-order correction. We discuss briefly in §6 the extension of co-sufficient model assessment to settings other than the Gaussian linear model.
Between the two values of $k$ reported, intermediate choices of $k$ interpolate between the reported values of Tables 45 in terms of coverage probability and the expected size of the confidence set, with small $k$ achieving higher coverage at the expense of larger confidence sets.
The strategy for reduction is not the focus of this paper. Nevertheless, the tables reflect aspects of the discussion of Lewis and Battey (2025). In particular, Cox reduction leads, on average, to slightly smaller confidence sets than the lasso for the same coverage, as the covariates retained by Cox reduction typically span a larger space than those retained by the lasso. As a result, the retained covariates from Cox reduction fit the data better, making it harder for small sub-models to pass the $F$-test. Cox reduction is also relatively better at retaining signal variables through the reduction phase when the correlation between signal and noise variables is high, and vice versa, as the lasso is more likely to select noise variables that are highly correlated with signal variables.
In all correctly calibrated cases, we see that the size of the confidence set of models is rather large. The maximal size of the confidence set obtained by assessing models up to size 5 from a reduced set with |^$|\hat{\mathcal{E}}|=15$ is 4943. Thus, all approaches have power to reject false models in these low-information settings. Cox and Battey (2017) discuss how to extract compact messages from large confidence sets of models.
@@ 6.1. Discussion and immediate extension
The ideas of §2.1 and §2.2 apply in principle to any parametric regression model in which, under the model hypothesis, a sufficient statistic or approximate sufficient statistic of smaller dimension than $n$ is available. An exact or approximate analytic approach for non-Gaussian models is likely to lead to more marked improvements over sample splitting relative to the Gaussian case, as the $F$ test for model adequacy is exact, conditional on the true model $\mathcal{E}^{*}$ having survived the first stage, while the log-likelihood test needed for more general models models is based on an asymptotic approximation.
@@
The extension to other models is less direct. In §6.2 and §6.4 we emphasise the structure in canonical exponential-family regression models and in accelerated life models that simplifies the required calculations. The logistic regression case is developed in a little more detail in §6.3, although without rigorous justification.
@@ 6.2. Canonical exponential family regression models
Provided that any floating dispersion parameter is known, the sufficient statistic for $\theta$ is $S=\sum_{i=1}^{n}x_{i}Y_{i}=X^{{\rm T}}Y$, where each $x_{i}$ is of the same dimension $d=d_{\theta}$ as $\theta$. There is a reduction by sufficiency from $n$ to $d$ and no ancillary statistic in the sense of §2.2. The insights of §6.5 to be presented show that the co-sufficient manifold $\mathcal{Q}(s^{o})$ is flat. We consider in more detail the example of logistic regression.
@@ 6.3. Logistic regression
In contrast to §3, where the conditional distribution of interest reduced to a marginal distribution by the properties of Gaussian distributions, the co-sufficient statistic $Q(s^{o})$ is notional in the sense that it cannot be expressed in terms of observable quantities and is given implicitly by the defining equation (2.3). The conditional distribution of interest is
$\mathop{\mathrm{pr}}(Y=y\mid S=s^{o})=\frac{\mathop{\mathrm{pr}}(Y=y)}{\sum_{y^{\prime}\in\mathcal{Q}(s^{o})}\mathop{\mathrm{pr}}(Y=y^{\prime})}.$ (6.1)
With a different objective in mind, Hanley and Roy (2025) obtained by linear integer programming, all sets of binary outcomes $y\in\{0,1\}^{n}$ compatible with the observed value of the sufficient statistic $s^{o}=X^{{\rm T}}y^{o}$, i.e. a complete enumeration of $\mathcal{Q}(s^{o})$, which is a finite set in this case. Their original motivation for obtaining these solutions was to re-analyse by a superficially more modern approach, an old analysis due to Penrose (1933), aided by Fisher. The conclusion of their pedagogical piece was that all solutions to this linear integer programming problem gave the same inference for the logistic regression parameters as that obtained by Penrose and Fisher, as they must by construction, by the definition of sufficiency. Summation over Hanley’s and Roy’s (2025) full solution set in equation (6.1) provides a means of calculating the conditional distribution of interest. In practice, the probabilities on the right hand side of (6.1) depend on the unknown parameter vector $\theta$, which can be replaced by the maximum likelihood solution $\hat{\theta}$.
@@ 6.4. Accelerated life models
$Y:=\log T=\log T_{0}-\log\rho(z)=\log T_{0}-z^{{\rm T}}\beta=\mu_{0}-z^{{\rm T}}\beta+\varepsilon,$ (6.2)
where $\mu_{0}=\mathbb{E}(\log T_{0})$ and $\varepsilon$ is a mean-zero error term, whose distribution is determined by that of the baseline variable $T_{0}$. The broad ideas of §3 apply to the log-transformed outcome model (6.2) with $\theta$ one dimension larger than $\beta$ so as to include the intercept $\mu_{0}$. If the baseline outcome distribution is log-normally distributed, no modification is needed. Here we consider an alternative choice of outcome distribution to highlight the considerations involved.
@@
$f_{\varepsilon}(\epsilon)=\frac{\exp(-\epsilon/\tau)}{\tau(1+\exp(-\epsilon/\tau))},$ (6.3)
@@
In view of the discussion of §4.5, we only consider inference based on $W$, rather than on $Q=U^{{\rm T}}Y/\|U^{{\rm T}}Y\|$. The latter has a distribution on the unit hypersphere in $\mathcal{X}^{\perp}$, however this distribution is not uniform and is more difficult to calculate in the present setting. The characteristic function $\varphi_{W}(z)$ of $W$ can be written in terms of the characteristic function $\varphi_{\varepsilon}$ of $\varepsilon$ as $\varphi_{W}(z)=\varphi_{\varepsilon}(U^{{\rm T}}z)$ where
$\varphi_{\varepsilon}(t)=\prod_{i=1}^{n}\varphi_{\varepsilon_{i}}(t_{i})=(\pi\tau)^{n}\prod_{i=1}^{n}\frac{t_{i}}{\sinh(\pi\tau t_{i})}.$ (6.4)
@@
[…]
where $\Re(\cdot)$ denotes the real part of a complex number. Voldoire et al. (2024) suggest a procedure for obtaining samples from (6.5). By comparing observed values of $W$ to the tails of this sampling-based distribution, adequacy of the model can be assessed. This requires the logistic standard deviation $\tau$ in (6.4) to be replaced by a consistent estimate, as in §4.3. Ideally, the integral (6.5) might instead be approximated analytically, as sampling-based approaches are prohibitively expensive in terms of computation for reasons mentioned earlier.
@@ 6.5. General non-exponential family models
$V(y^{o})=\left.\frac{\partial s^{{\rm T}}(y)}{\partial y}\right|_{y=y^{o}}$ (6.6)
specify the normal directions to $\mathcal{Q}(s^{o})$ at $y^{o}$. The space spanned by the columns of $V(y^{o})$ is the normal subspace of $\mathcal{Q}(s^{o})$ at $y^{o}$ and the orthogonal complement
@@
defines the tangent plane at $y^{o}$. For canonical exponential-family regression models with no floating dispersion parameter these matrices do not depend for $y^{o}$, implying that the co-sufficient manifold is flat. This idea underpins the tangent exponential model of Fraser (1988, 1990) and Fraser and Reid (1988, 1995), whose objective was higher order conditional inference on $\theta$. Our problem is in essence the opposite of theirs. The establishment of general analytic approximations is an open problem. In principle, such approximations could replace the importance-sampling based approach suggested for the log-logistic accelerated life model of the previous section, and possibly also the integer-programming based approach for the logistic regression formulation of §6.3.
@@ 7. Closing discussion
The literature on post-selection inference has focused on the conceptually different problem of performing inference on a parameter following a data-based selection of effects on which formal inference is sought. The problem of post-reduction inference for confidence sets of models raises similar challenges, but is conceptually very different. In particular, the object of inference in the confidence sets of models problem is unambiguously defined, and the two phases of analysis are such that there is a clearer separation of information in the data for the different phases. This contrasts strongly with the more widely studied post-selection inference problems, which address almost the same scientific question in different guises: a selection phase asking whether the effect is zero, and an inference phase asking what values of strongly-suggested effects are compatible with the data.
In the confidence sets of models formulation, the most relevant hypotheses for test, as small submodels of $\hat{\mathcal{E}}$, are stable over hypothetical replication from the same generating process. Provided that $\hat{\mathcal{E}}$ reliably includes $\mathcal{E}^{*}$, the double use of the data is only problematic if the overfitted $\hat{\mathcal{E}}$ is used as a basis for comparison of fit, as in the likelihood ratio test. The present paper showed how to avoid this comparison by operationalising the appropriate co-sufficiency and ancillary separations. If the ancillary statistic for the interest component depends weakly on nuisance parameters as in the weaker definitions of ancillarity on page 38 of Barndorff-Nielsen and Cox, then the small-sample coverage properties of the ancillary-based assessment may be compromised through estimation of these nuisance parameters. Numerical work showed that, although the efficiency of sample splitting is relatively high for the Gaussian model under consideration here due to the exactness of the second-stage test, there are modest advantages of co-sufficient information partitioning in settings where the reduction phase is difficult, for instance when the sample size is small relative to dimension or the signal strength is low.
@@
A natural first reaction to the tables in §5 is that the number of models in the confidence set is inconveniently large. This is an honest reflection of the limitations of the data, and emphasises the dangers of reporting a single model when many are compatible with the data at the chosen significance level. Statistics can take us no further, and any choice between models in the confidence set requires either additional data or subject-matter expertise. Cox and Battey (2017) discuss how one might extract compact messages from large confidence sets of models. Specifically, among the large number of models reported in their supplementary material, two variables, call them $v_{1}$ and $v_{2}$, are present in 96% and 94% of cases. In 78% of the models in which $v_{2}$ is not present, another variable, $v_{3}$, is present in its place, and only 1% of models include neither $v_{2}$ nor $v_{3}$.
If prediction is the sole objective, an arbitrary well-fitting model, such as that selected by the lasso, will be equally effective and there is no need for a confidence set of models, although they could be used to provide a confidence distribution for predictions as discussed in Section S3 of the supplement of Lewis and Battey (2025).
@@ Acknowledgements
HB is grateful to James Hanley for sharing a preprint of Hanley and Roy (2025). EPSRC fellowship EP/T01864X/1 is gratefully acknowledged.
@@
Appendix A Proofs and derivations
A.1. Derivation of equation (4.2)

### v2

@@ 6.2. Coverage and expected size of the confidence sets
Simulations (not reported) with the distribution of the error term taken as student-$t$ gave qualitatively similar conclusions. In that case, the distribution theory for the oracle version of the co-sufficient test remains valid, but that for the $F$ test is invalidated. The survival probabilities of the full set of signal variables are lowered in this case, and the distribution theory for $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ established in Proposition 5.2 is no longer valid, affecting Theorem 5.3.
In the Gaussian context, because the $F$ test is exact, the coverage probability is only affected by the reduced survival probability in the first stage and not by the small sample size at the second stage, and the efficacy of sample splitting is high, the directionality of the $F$ test compensating to a large extent for the smaller sample size, provided that all signal variables survive reduction. In broad terms, therefore, the main aspect influencing the relative advantage of the co-sufficient and ancillary tests over the split-sample tests is the difficulty of the reduction phase. In non-linear regression models, where a likelihood ratio test has to be used in place of the $F$ test, the coverage probability of $\mathcal{M}$ is also affected by the smaller sample size in the second stage, due to the $\chi^{2}$ approximation to the likelihood-ratio statistic, although this too could be improved by higher-order correction. We discuss briefly in §7 the extension of co-sufficient model assessment to settings other than the Gaussian linear model.
Between the two values of $k$ reported, intermediate choices of $k$ interpolate between the reported values of Tables 1–5 in terms of coverage probability and the expected size of the confidence set, with small $k$ achieving higher coverage at the expense of larger confidence sets.
The strategy for reduction is not the focus of this paper. Nevertheless, the tables reflect aspects of the discussion of Lewis and Battey (2025). In particular, Cox reduction leads, on average, to smaller confidence sets than the lasso for the same coverage, as the covariates retained by Cox reduction typically span a larger space than those retained by the lasso. As a result, the retained covariates from Cox reduction fit the data better, making it harder for small sub-models to pass the $F$-test. Cox reduction is also relatively better at retaining signal variables through the reduction phase when the correlation between signal and noise variables is high, and vice versa, as the lasso is more likely to select noise variables that are highly correlated with signal variables.
In all correctly calibrated cases, we see that the size of the confidence set of models is rather large, particularly for the co-sufficient test. Small errors of calibration seem worth the appreciable reduction in the number of models retained from splitting the sample. The large errors of coverage, as reported in Table 1, are more severe, making the co-sufficient approach more relevant at small sample sizes. The maximal size of the confidence set obtained by assessing models up to size 5 from a reduced set with $|\hat{\mathcal{E}}|=15$ is 4943. Thus, all approaches have power to reject false models in these low-information settings. Cox and Battey (2017) discuss how to extract compact messages from large confidence sets of models.
@@ 7.1. Discussion and immediate extension
The ideas of §3.2 and §3.3 extend in principle to any parametric regression model admitting a sufficient, or approximately sufficient, statistic of dimension smaller than $n$. An exact or approximate analytic approach for non-Gaussian models is likely to lead to more marked improvements over sample splitting relative to the Gaussian case, as the $F$ test for model adequacy is exact, conditional on the true model $\mathcal{E}^{*}$ having survived the first stage, while the log-likelihood test needed for more general models is based on an asymptotic approximation.
@@
The extension to other models is less direct and it remains unclear whether a randomisation approach paralleling §5.2 would be needed, as its necessity in the Gaussian case arose from rotational invariance of the Haar measure. Since the notional co-sufficient statistic is not uniformly distributed in any of the examples of the present section, a degree of power would be expected without any synthetic replication. Further work is needed to understand these cases.
In §7.2 and §7.4 we emphasise the structure in canonical exponential-family regression models and in accelerated life models that simplifies the required calculations.
@@ 7.2. Canonical exponential family regression models
Provided that any floating dispersion parameter is known, the sufficient statistic for $\theta$ is $S=\sum_{i=1}^{n}x_{i}Y_{i}=X^{{\rm T}}Y$, where each $x_{i}$ is of the same dimension $d=d_{\theta}$ as $\theta$. There is a reduction by sufficiency from $n$ to $d$ and no ancillary statistic in the sense of §3.3. The insights of §7.5 show that the co-sufficient manifold $\mathcal{Q}(s^{o})$ is flat. We consider in more detail the example of logistic regression.
@@ 7.3. Logistic regression
Unlike the Gaussian case of §4, where the conditional distribution reduces to a marginal distribution, the co-sufficient statistic $Q(s^{o})$ in logistic regression is not observable and is defined only implicitly through (3.3). The conditional distribution of interest is
(7.1) $\mathop{\mathrm{pr}}(Y=y\mid S=s^{o})=\frac{\mathop{\mathrm{pr}}(Y=y)}{\sum_{y^{\prime}\in\mathcal{Q}(s^{o})}\mathop{\mathrm{pr}}(Y=y^{\prime})}.$
With a different objective in mind, Hanley and Roy (2025) obtained, by linear integer programming, all sets of binary outcomes $y\in\{0,1\}^{n}$ compatible with the observed value of the sufficient statistic $s^{o}=X^{{\rm T}}y^{o}$, i.e. a complete enumeration of $\mathcal{Q}(s^{o})$, which is a finite set in this case. Their original motivation for obtaining these solutions was to re-analyse by a superficially more modern approach, an old analysis due to Penrose (1933), aided by Fisher. The conclusion of their pedagogical piece was that all solutions to this linear integer programming problem gave the same inference for the logistic regression parameters as that obtained by Penrose and Fisher, as they must by construction, by the definition of sufficiency. Summation over Hanley’s and Roy’s (2025) full solution set in equation (7.1) provides a means of calculating the conditional distribution of interest. In practice, the probabilities on the right hand side of (7.1) depend on the unknown parameter vector $\theta$, which can be replaced by the maximum likelihood solution $\hat{\theta}$.
@@ 7.4. Accelerated life models
(7.2) $Y:=\log T=\log T_{0}-\log\rho(z)=\log T_{0}-z^{{\rm T}}\beta=\mu_{0}-z^{{\rm T}}\beta+\varepsilon,$
where $\mu_{0}=\mathbb{E}(\log T_{0})$ and $\varepsilon$ is a mean-zero error term, whose distribution is determined by that of the baseline variable $T_{0}$. The broad ideas of §4 apply to the log-transformed outcome model (7.2) with $\theta$ one dimension larger than $\beta$ to incorporate the intercept $\mu_{0}$. If the baseline outcome distribution is log-normally distributed, no modification is needed. Here we consider an alternative choice of outcome distribution to highlight the considerations involved.
@@
(7.3) $f_{\varepsilon}(\epsilon)=\frac{\exp(-\epsilon/\tau)}{\tau(1+\exp(-\epsilon/\tau))},$
@@
We only consider inference based on $W$, rather than on $Q=U^{{\rm T}}Y/\|U^{{\rm T}}Y\|$. The latter has a distribution on the unit hypersphere in $\mathcal{X}^{\perp}$, however this distribution is not uniform and is more difficult to calculate in the present setting. The characteristic function $\varphi_{W}(z)$ of $W$ can be written in terms of the characteristic function $\varphi_{\varepsilon}$ of $\varepsilon$ as $\varphi_{W}(z)=\varphi_{\varepsilon}(U^{{\rm T}}z)$ where
(7.4) $\varphi_{\varepsilon}(t)=\prod_{i=1}^{n}\varphi_{\varepsilon_{i}}(t_{i})=(\pi\tau)^{n}\prod_{i=1}^{n}\frac{t_{i}}{\sinh(\pi\tau t_{i})}.$
@@
(7.5) $f_{W}(w)=\frac{1}{(2\pi)^{(n-d_{\theta})}}\int_{[-\pi,\pi]}\cdots\int_{[-\pi,\pi]}\Re\{\exp(-iz^{{\rm T}}w)\varphi_{\varepsilon}(U^{{\rm T}}z)\}dz_{1}\cdots dz_{n-d_{\theta}},$
where $\Re(\cdot)$ denotes the real part of a complex number. Voldoire et al. (2024) suggest a procedure for obtaining samples from (7.5). By comparing observed values of $W$ to the tails of this sampling-based distribution, adequacy of the model can be assessed. This requires the logistic standard deviation $\tau$ in (7.4) to be replaced by a consistent estimate, as in §5.2. Ideally, (7.5) would admit a tractable analytic approximation, as direct sampling is computationally prohibitive for the reasons noted earlier.
@@ 7.5. General non-exponential family models
(7.6) $V(y^{o})=\left.\frac{\partial s^{{\rm T}}(y)}{\partial y}\right|_{y=y^{o}}$
specifies the normal directions to $\mathcal{Q}(s^{o})$ at $y^{o}$. The space spanned by the columns of $V(y^{o})$ is the normal subspace of $\mathcal{Q}(s^{o})$ at $y^{o}$ and the orthogonal complement
@@
defines the tangent plane at $y^{o}$. For canonical exponential-family regression models without a floating dispersion parameter, $V(y^{o})$ does not depend on $y^{o}$, implying that the co-sufficient manifold is flat. This idea underpins the tangent exponential model of Fraser (1988, 1990) and Fraser and Reid (1988, 1995), whose objective was higher order conditional inference on $\theta$. Our problem is in essence the opposite of theirs. The establishment of general analytic approximations is an open problem. In principle, such approximations could replace the importance-sampling based approach suggested for the log-logistic accelerated life model of the previous section, and possibly also the integer-programming based approach for the logistic regression formulation of §7.3.
@@ 8. Closing discussion
The literature on post-selection inference has focused on the conceptually different problem of performing inference on a parameter following a data-based selection of effects on which formal inference is sought. The problem of post-reduction inference for confidence sets of models raises similar technical challenges, but is conceptually distinct. In particular, the object of inference in the confidence sets of models problem is unambiguously defined, and the two phases of analysis are such that there is a clearer separation of information in the data for the different phases. This contrasts strongly with the more widely studied post-selection inference problems, which address almost the same scientific question in different guises: a selection phase asking whether the effect is zero, and an inference phase asking what values of strongly-suggested effects are compatible with the data.
In the confidence sets of models formulation, the most relevant hypotheses for test, as small submodels of $\hat{\mathcal{E}}$, are stable over hypothetical replication from the same generating process. Provided that $\hat{\mathcal{E}}$ reliably includes $\mathcal{E}^{*}$, the double use of the data is only problematic if the overfitted $\hat{\mathcal{E}}$ is used as a basis for comparison of fit, as in the likelihood ratio test. The present paper showed how to avoid this comparison by operationalising the appropriate co-sufficiency and ancillary separations. If the ancillary statistic for the interest component depends weakly on nuisance parameters as in the weaker definitions of ancillarity on page 38 of Barndorff-Nielsen and Cox (1994), then the small-sample coverage properties of the ancillary-based assessment may be compromised through estimation of these nuisance parameters. Numerical work showed that, although the efficiency of sample splitting is relatively high for the Gaussian model under consideration here due to the exactness of the second-stage test, there are modest advantages of co-sufficient information partitioning in settings where the reduction phase is difficult, for instance when the sample size is small relative to dimension or the signal strength is low.
@@
A natural first reaction to the tables in §6 is that the number of models in the confidence set is inconveniently large. This is an honest reflection of the limitations of the data, and emphasises the dangers of reporting a single model when many are compatible with the data at the chosen significance level. Statistical analysis can go no further, and any choice between models in the confidence set requires either additional data or subject-matter expertise. Cox and Battey (2017) discuss how one might extract compact messages from large confidence sets of models. Specifically, among the large number of models reported in their supplementary material, two variables, call them $v_{1}$ and $v_{2}$, are present in 96% and 94% of cases. In 78% of the models in which $v_{2}$ is not present, another variable, $v_{3}$, is present in its place, and only 1% of models include neither $v_{2}$ nor $v_{3}$.
If prediction is the sole objective, an arbitrary well-fitting model, such as that selected by the lasso, will typically suffice, and a confidence set of models is unnecessary. Confidence sets may nevertheless be used to form confidence distributions for predictions, as discussed in Section S3 of the supplement of Lewis and Battey (2025).
@@ Acknowledgements
HB is grateful to James Hanley for sharing a preprint of Hanley and Roy (2025). We thank two reviewers and the associate editor for detailed reading, thoughtful suggestions, and for indicating an oversight in the proof of Proposition 2. EPSRC fellowship EP/T01864X/1 (to HB) is gratefully acknowledged.
@@
The supplementary material contains Appendices A and B referenced in the main text. These include proofs and technical details.
SUPPLEMENTARY MATERIAL
Appendix A Power and naïve data-partitioning
From $k$ disjoint subsamples of size $n_{k}=n/k$, we consider initially $Q^{(1)},\ldots,Q^{(k)}$, constructed as above but based on disjoint subsamples $(Y^{(1)},X^{(1)}),\ldots,(Y^{(k)},X^{(k)})$.
(A.1) $\frac{\exp\{-(a+w^{{\rm T}}w)/2\sigma^{2}\}}{(2\pi\sigma^{2})^{(n-d_{\theta})/2}}\exp\Bigl(-\frac{w^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma^{2}}\Bigr)dw,$
In the absence of a closed form for the marginal distribution of $Q$, Appendix §B.1 establishes the approximation
(A.2) $f_{Q}(q)d\text{Vol}^{n-d}(q)\simeq\frac{\Gamma\bigl(\frac{n-d_{\theta}}{2}\bigr)}{2\pi^{(n-d_{\theta})/2}}\exp\Bigl(-\frac{a}{2\sigma^{2}}\Bigr)\exp\biggl\{\frac{(n-d_{\theta})q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma\sqrt{2}}\biggr\}d\text{Vol}^{n-d}(q),$
which we write as
Approximation (A.2), although obtained under the notional limiting operation $(n-d_{\theta})\rightarrow\infty$, is essentially exact, the only error coming from Stirling’s approximation.
The first multiplicative factor on the right hand side of (A.2) is the density function of $Q$ under the null hypothesis. Thus, power is low when $h(q)$ is close to 1 for all $q$. This occurs when $Z\lambda$ is close to zero or when $Z$ lies close to the column span of $X$. The conclusion agrees qualitatively with the power analysis for the likelihood-ratio test in the absence of any dependencies from reduction, as derived in §4.2 of Lewis and Battey (2025) using a limiting argument in Le Cam-type contiguous alternatives.
Consider now a test of uniformity on the unit hypersphere in $\mathbb{R}^{n_{k}-d_{\theta}}$ based on the partitioned data, discarding the orientation of each $\mathcal{X}^{(i)\perp}$ within $\mathbb{R}^{n_{k}}$; $k$ now plays the role of sample size for the null uniform samples. Equation (A.2) suggests that the non-null distribution of $Q^{(i)}$ is well approximated by a von Mises-Fisher distribution with mean direction $m^{(i)}=U^{(i)T}Z^{(i)}\lambda$ and dispersion parameter $\kappa^{(i)}=\sigma^{-1}2^{-1/2}(n_{k}-d_{\theta})\|U^{(i)T}Z^{(i)}\lambda\|$. Since the mean vectors $m^{(i)}$ depend on how the data are partitioned, and because these themselves are in general numerically indistinguishable from uniformly distributed random vectors on the unit hypersphere in $\mathbb{R}^{n_{k}-d_{\theta}}$ when standardised to unit length, the distribution of the test statistic when the assumed model is violated is hard to distinguish from that under the model postulated under the null hypothesis. Put differently, by ignoring the orientation of the embedding manifold $\mathcal{X}^{(i)\perp}$, the analysis is made too unconditional.
Appendix B Proofs and derivations
B.1. Derivation of equation (A.2)